# Novel point cloud registration approach for noninvasive patient specific estimation of leaflet strain from 3D images of heart valves

### v1

@@ 3 Results
We performed comprehensive leaflet strain analyses on three patient cohorts from the Children’s Hospital of Philadelphia and the University of Pennsylvania. Detailed verification results on shape agreement using our proposed method, together with a comparative analysis against other point-based registration techniques, are provided in Section 3.1. Section 3.2 presents strain analysis of TVs in patients with hypoplastic left heart syndrome (HLHS). Section 3.3 provides strain analysis of pediatric MVs, including both normal and a range of pathologic conditions, such as rheumatic mitral regurgitant valve and Marfan syndrome. Section 3.4 covers the strain analysis of normal adult MVs. In this context, “normal” indicates valves that coapt properly without regurgitation or tissue calcification. However, these adult patients had other cardiac conditions, such as aortic stenosis or coronary artery disease, at the time of imaging. All registration-based strain analyses were performed on a MacBook Pro with an Apple M2 Max chip (Apple Computer, Cupertino, CA). The average computational time for registration and strain estimation combined was 14.2 seconds per valve (with 3000 nodes per valve).
3.1 Verification results
@@
An image-derived mitral valve (MV) model from Wu et al.[40] was used as the benchmark to verify the proposed algorithm. The ground truth valve deformation, areal strain, and 1 principal Green Lagrange strain were generated in FEBio using the same model setup described in the original study. The strain formulation is described in Section 2.3. We compared the registration and strain estimation performance of our method against established point-based registration approaches: Bayesian-CPD (BCPD) [41], support vector registration (SVR) [42], and traditional CPD [26]. Probreg [43] was used to perform the BCPD and SVR registration.
The reference areal and 1 principal strain fields are shown in Fig. 2A. Fig. 2B provides a qualitative comparison between the estimated and ground truth strain maps. As illustrated, both BCPD and SVR failed to register the deformed geometry of the MV. In contrast, the standard CPD and our proposed DW-CPD successfully captured both the deformed shape and the associated strain patterns. The absolute difference between registration-informed strain and the FEA solution from CPD and DW-CPD approach is plotted in Fig. 2C. We observed higher absolute errors in both areal and first principal strain in regions with high curvature and bending (e.g., near the commissure folds) for both methods. CPD overpredicted strain more than DW-CPD, with up to 0.8 absolute error in area strain. Quantitative comparisons reporting errors in shape agreement and strain estimation are summarized in Fig. 2D. Shape agreement was assessed using the mean symmetric distance (MSD) and the 95 %ile Hausdorff distance (HD), while strain accuracy was evaluated using the mean absolute error (MAE) of the areal and 1 principal strains. Across all four metrics, our method consistently outperformed others, achieving an MSD of 0.32 mm, a 95 %ile HD of 0.71 mm, an areal strain MAE of 0.15, and a 1 principal strain MAE of 0.07.
@@
We applied the proposed DW-CPD feature-tracking method to estimate leaflet deformation in three patient cohorts: (1) TV in patients with HLHS, (2) pediatric MVs, and (3) adult MVs. Fig. 3A shows an example where the CPD method produced a registration artifact at the leaflet edge, causing it to fold inward and protrude, resulting in self-penetration. In contrast, DW-CPD accurately tracked the leaflet deformation, eliminating this defect. The mean and SD of the MSD and 95 percentile HD for each leaflet within each cohort are summarized in Fig. 3B. DW-CPD was used to inform the strain calculation among the patient cohorts. Overall, the registered leaflets achieved strong agreement compared with the ground truth segmentation models. Among the patient groups, the anterior leaflet of HLHS TVs showed the largest errors, with an average MSD of 0.357mm 0.068 and a 95 %ile HD of 0.857 mm 0.157.
3.2 Strain assessment of pathological tricuspid valves
@@ 3.2.2 Strain characteristics on individual leaflet
We compared the average areal and 1 principal strain in the trivial to mild TR and mild or greater TR cohorts. The mean and standard deviation (SD) of the average areal strain and 1 principal strain on each leaflet are shown in Fig. 4D. The average areal strain quantifies the relative change in leaflet surface area between diastolic and systolic frames, whereas the average 1 principal strain represents the mean maximum tensile strain across the leaflet. With respect to areal strain, both groups exhibited negligible changes in septal leaflet area while demonstrating increased anterior and posterior surface area. Both the trivial to mild TR and mild to greater TR cohorts showed similar anterior areal strain, although the trivial to mild TR had slightly lower values (0.025 0.120 vs. 0.075 0.124). However, posterior areal strain was substantially higher in the trivial to mild TR cohort compared to the mild or greater cohort (0.302 0.381 vs. 0.106 0.282). This suggests that patients in the trivial and mild cohort have more extensible posterior leaflets. Consistent with the areal strain findings, the trivial to mild TR cohort exhibited lower average 1 principal strain in the anterior leaflet compared to the mild or greater cohort (0.303 0.113 vs. 0.410 0.158), but higher average 1 principal strain in the posterior leaflet (0.517 0.221 vs. 0.463 0.203). For the septal leaflet, the average 1 principal strain was 0.273 0.026 in the trivial to mild cohort and 0.361 0.206 in the mild or greater cohort.
@@ 3.2.3 Population study of strain patterns across the patient cohort
Principal component analyses (PCA) of the TVs at mid-diastole and mid-systole were performed independently to determine the mean valve shape and its standard deviations (SD) across the 11 patients. Leaflet strains were computed by tracking point-to-point correspondences between the open and closed valve geometries using our distance-weighted coherent point drift (DW-CPD) registration approach. The resulting spectrum of valve morphologies and associated strain information is shown in Fig. 4E. Within this spectrum, the valve appeared most tethered at –2 SD from the mean shape and most billowing at +2 SD. Both ends of the morphological spectrum demonstrated markedly elevated strains relative to the mean shape. In the most tethered valve, high strain was distributed across all three leaflets, whereas in the most billowing valve, high strain was concentrated on the septal leaflet.
The mean shape exhibited relatively low strains (–0.114 0.259 average areal strain and 0.148 0.175 average 1 principal strain). In contrast, the most tethered valve showed an average areal strain of 0.169 0.433 and a 1 principal strain of 0.425 0.246, while the most billowing valve demonstrated an average areal strain of –0.203 0.392 and a 1 principal strain of 0.221 0.300. For both areal and 1 principal strain, the SD of the strain distribution increased as valve geometries deviated further from the mean shape. These findings suggest that both elevated 1 principal strain and greater strain variability, represented by SD, are associated with pathological TV morphology.
3.3 Strain assessment of normal and diseased pediatric mitral valves
@@ 3.3.1 Patient population
This study cohort included 10 pediatric patients, aged 2 to 16 years at the time of image acquisition. Of these, 5 had normal coaptation, 3 demonstrated billowing of one or more leaflets, and 2 had restricted posterior leaflets. Further, the cohort included 1 patient with an MV arcade, 1 with rheumatic MV, and 1 with Marfan syndrome. Fig. 5A shows the superimposition of ground truth segmentation and the estimated atrial leaflet surface onto 3D TEE images from two representative patients. The ages and clinical diagnoses of all patients are summarized in Fig. 5B.
@@ 3.3.2 Strain characteristics on individual leaflet
We organized the patient cohort into normal and diseased groups and compared the average areal and 1 principal strains (Fig. 5D). Neither group exhibited significant changes in leaflet area between the mid-diastolic and mid-systolic frames. In the normal cohort, the average areal strain was 0.034 0.147 in the anterior leaflet and 0.029 0.171 in the posterior leaflet. In the diseased cohort, the corresponding values were –0.038 0.198 and 0.018 0.354.
The posterior leaflets exhibited higher 1 principal strain than the anterior leaflets in both normal and diseased cohorts. In the normal cohort, the average 1 principal strain was 0.262 0.152 in the anterior leaflet and 0.337 0.117 in the posterior leaflet. In the diseased cohort, the corresponding values were 0.223 0.124 and 0.388 0.356. No clear differences in 1 principal strain were observed between the two cohorts, as the inter-cohort differences for each leaflet were less than 0.05.
@@ 3.3.3 Population study of strain patterns across the patient cohort
Although differences in MV strain were modest, we observed a unidirectional increase in mean areal strain from –0.084 to –0.035 as MV morphology shifted from –2 to +2 SD of valve shapes. In contrast, both ends of the morphological spectrum exhibited higher average 1 principal strain than the mean shape: 0.162 at –2 SD and 0.191 at +2 SD, compared to 0.131 in the mean shape. Similar to the TV analysis, the SD of areal and 1 principal strain increased as valve geometries deviated further from the mean.
3.4 Strain assessment of adult mitral valves
@@ 3.4.1 Patient population
This study cohort comprised 10 adult cardiac surgery patients, aged 49 to 78 years at the time of intra-operative image acquisition. All patients demonstrated normal MV function with no to mild mitral regurgitation (MR), though each had non-mitral cardiac pathology. The ground truth segmentation and the estimated atrial leaflet surface overlaid onto 3D TEE images from two representative patients are presented in Fig.6A. Further, Fig.6B summarizes the ages and clinical diagnoses of the entire cohort. Unlike the previous two cohorts, none of the patients in this group were identified with billowing leaflets.
@@ 4.1 General comments
In this study, we present an open-source, registration-based framework for quantifying leaflet strain in atrioventricular valves. We verified our implementations against FEA benchmark, and excellent shape and strain agreements were achieved. Compared with BCPD [41], SVR [42], and traditional CPD [26], our method consistently achieved superior performance across all metrics. To assess robustness to morphological variation, we applied the framework to three patient cohorts: (1) TVs in patients with HLHS, (2) pediatric MVs, and (3) adult MVs. The dataset spans normal to regurgitant valves and ages from neonates to the late seventies. This constitutes the first human analysis of tricuspid valve strain quantification comparing healthy and diseased cohorts, and the most heterogeneous patient population studied to date. Despite substantial variation in valve geometry across patients and cohorts, the same registration parameters produced accurate results in all cases. This highlights the generalizability of the framework. Across cohorts, the median and IQR of the 1 principal strain distributions emerged as the most sensitive biomechanical marker of leaflet health, with elevated median and IQR values of 0.5 consistently indicating billowing.
The assessment of the biomechanical strain of valves is central to the understanding of the mechanical environment and resulting tissue and cellular changes underlying disease initiation and progression [10, 11, 12, 13]. FEA has been widely used to examine the effects of annular dynamics [19, 20], papillary muscle displacement [44], constitutive properties [45, 46], and annular geometry [47] on valve function. Yet most studies rely on animal or in vitro models with parameters derived from non-human tissue. Although image-derived FEA now allows patient-specific reconstructions from clinical imaging [48, 16, 21], challenges remain. Fitting of FEM based structures to image derived valve leaflet models can be challenging due to complex leaflet morphology. In addition, such approaches can be prone to user bias as a result of the need for significant local parameter tuning to reliably capture the leaflet surface with high fidelity. Further, such simulations require assumptions related to subvalvar chordal attachments or a surrogate for these forces [16, 21]. In contrast, registration-based strain offers a compelling alternative for quantifying strain from 3DE, as the formulation does not require assumptions about constitutive models and can be performed on the leaflet surfaces themselves. While registration-based approaches have been explored, application to valvular strain has been limited to mitral valve assessment with dependence on input from commercial mitral valve modeling software [15]. Further, this existing capability tracked valves from mid-systole to end-systole, which does not capture the near strain free diastolic state of an open valve as a baseline. In addition, this existing software cannot be extended to accommodate diverse valves nor allow incorporation of segmentation-based valve models, which can increasingly be rapidly generated using machine learning-based approaches [49, 50]. Finally, while 3D + time (4D) CT and cardiac magnetic resonance imaging are increasingly being utilized to evaluate valve function, there is currently no generalizable framework for analysis of valve leaflet strain across any 3D modality with sufficient temporal resolution. The need for multimodality-based assessment of valve leaflet strain across diverse valve morphologies and populations, including patients with complex congenital heart disease, is the key motivation for the present study.
@@
To provide a summary of our strain analysis for MVs, clear differences in the characteristics of 1 principal strain distributions between pediatric and adult MVs were observed. In pediatric valves, the distribution was broader with a higher IQR, whereas in adults, it was narrower and more peaked. In population-level PCA, both groups showed increasing strain IQR as valve geometries deviated from the mean shape. In adults, distinct patterns emerged in which strain in the posterior leaflet progressively increased and became more diffuse with increasing anterior–posterior diameter. In contrast, this pattern was not observed in the pediatric MV cohort. In our pediatric cohort (mean age 8.9 4.9 years), the average principal strain was 0.262 in the anterior leaflet vs. 0.337 in the posterior leaflet in valves with normal morphology. Conversely, the average 1 principal strain reported in our adult cohort (mean age 64.6 9.4 years) was markely lower, with strains exhibiting 0.081 and 0.231, respectively. These findings confirm the general intuition that leaflet strain reduces with age, attributed to an increase in tissue stiffness and thickness.

### v2

@@ 3 Results
We performed comprehensive leaflet strain analyses on three patient cohorts from the Children’s Hospital of Philadelphia and the University of Pennsylvania. Detailed verification results on shape agreement using our proposed method, together with a comparative analysis against other point-based feature tracking techniques, are provided in Section 3.1. Section 3.4 presents strain analysis of TVs in patients with hypoplastic left heart syndrome (HLHS). Section 3.5 provides strain analysis of pediatric MVs, including both normal and a range of pathologic conditions, such as rheumatic mitral regurgitant valve and Marfan syndrome. Section 3.6 covers the strain analysis of normal adult MVs. In this context, “normal” indicates valves that coapt properly without regurgitation or tissue calcification. However, these adult patients had other cardiac conditions, such as aortic stenosis or coronary artery disease, at the time of imaging. All feature tracking-based strain analyses were performed on a MacBook Pro with an Apple M2 Max chip (Apple Computer, Cupertino, CA). The average computational time for feature tracking and strain estimation combined was 14.2 seconds per HLHS tricuspid valve, using meshes with 1000 nodes per leaflet.
3.1 Finite element analysis verification
@@
An image-derived mitral valve (MV) model from Wu et al.[51] was used as the benchmark to verify the proposed algorithm. The ground truth valve deformation, areal strain, and 1 principal Green Lagrange strain were generated in FEBio using the same constitutive model parameters and pressure. The undeformed (mid-diastolic) mesh and the deformed (mid-systolic) mesh were used as inputs to all feature tracking methods to evaluate their performance in approximating the deformed geometry and leaflet surface strain relative to FE–derived approximations. The strain formulation is described in Section 2.3. We compared the feature tracking and strain estimation performance of DW-CPD against established point-based feature tracking approaches: Bayesian-CPD (BCPD) [14] and classic CPD [28]. Probreg [17] was used to facilitate the BCPD feature tracking.
The reference areal and 1 principal strain are shown in Fig. 2A. Fig. 2B provides a qualitative comparison between the estimated and ground truth strain maps. As illustrated, although all feature tracking methods capture the deformed systolic shape to a large extent, noticeable differences are observed in the estimated strain distributions across the three methods. BCPD substantially underpredicts strain on the anterior leaflet of the mitral valve compared with CPD and DW-CPD. Both CPD and DW-CPD capture strain across the entire valve. However, CPD tends to overpredict strain in regions experiencing high tensile bending. Shape agreement was assessed using the MSD and HD95, while strain accuracy was evaluated using the mean absolute error (MAE) of areal strain and first principal strain. Differences in the estimated systolic valve surface geometry and leaflet strain distributions are shown in Fig. 2C. Both CPD and DW-CPD achieved substantially lower MSD and HD95 than BCPD, with more than a three-fold improvement in shape similarity, and DW-CPD yielded the lowest geometric error. In addition, DW-CPD produced an areal strain MAE of 0.138 and a 1 principal strain MAE of 0.055, the lowest among all methods evaluated.
Visualizations of differences in areal strain and 1 principal strains between DW-CPD and FEA are shown in Fig. 2D. The DW-CPD appears to overpredict compressive and tensile strain in high curvature regions, especially near commissure folds. Fig. 2E presents scatter plots of areal strain and principal strain estimated by DW-CPD and FEA. The results demonstrate strong agreement between the two methods. For areal strain, the correlation was high, with Pearson’s () and Spearman’s (). For principal strain, Pearson’s was () and Spearman’s was ().
3.2 Segmental leaflet atrial surface evaluation
Table 1 summarizes the surface area of the leaflets during mid-diastole and mid-systole for each segmental atrial surface across the three patient cohorts, along with the corresponding percent differences in surface area between the two phases. The differences between diastolic and systolic surface areas varied, showing both increases and decreases across leaflets and patients. This resulted in a wide distribution of percent differences. In the HLHS TVs, leaflet surface area changes were mixed, with several showing positive percent differences while others had values close to zero or negative. A similar variability was noted in the pediatric MVs. In the adult MVs, negative percent differences were more frequently observed in both the anterior and posterior leaflets, although some positive changes in surface area were still present.
3.3 Quantitative assessment of feature-tracking accuracy
@@
We applied the proposed DW-CPD feature-tracking method to estimate leaflet deformation in three patient cohorts: (1) TV in patients with HLHS, (2) pediatric MVs, and (3) adult MVs. Fig. 3A shows an example where the CPD method produced a feature tracking artifact at the leaflet edge, causing it to fold inward and protrude, resulting in self-penetration. In contrast, DW-CPD accurately tracked the leaflet deformation, eliminating this defect. The mean and SD of the MSD and HD95 for each leaflet within each cohort are summarized in Fig. 3B. DW-CPD was used to inform the strain calculation among the patient cohorts. Overall, the registered leaflets achieved strong agreement compared with the ground truth segmentation models. Among the patient groups, the anterior leaflet of HLHS TVs showed the largest errors, with an average MSD of 0.357mm 0.068 and HD95 of 0.857 mm 0.157. While our primary goal was to demonstrate that our methods could capture leaflet geometry across diverse valve types and morphologies, and we were not powered to determine granular associations of strain to morphology, we explored descriptive statistics of leaflet geometry and leaflet strain across all three patient populations as described in the following section. This serves as a demonstration of how our methods could be applied to investigate associations of strain to valve disease and morphology in future studies, and to provide initial insights in pediatric patients where understanding of the relationships between valve morphology, strain, and pathology is nascent.
3.4 Strain assessment of pathological tricuspid valves (N=11)
@@ 3.4.2 Strain characteristics on individual leaflet
We compared the average areal and 1 principal strain in the trivial to mild TR and mild or greater TR cohorts. The mean and standard deviation (SD) of the average areal strain and 1 principal strain on each leaflet are shown in Fig. 4D. The average areal strain quantifies the relative change in leaflet surface area between diastolic and systolic frames, whereas the average 1 principal strain represents the mean maximum tensile strain across the leaflet. With respect to areal strain, both groups exhibited negligible changes in septal leaflet area while demonstrating increased anterior and posterior surface area. Both the trivial to mild TR and mild to greater TR cohorts showed similar anterior areal strain, although the trivial to mild TR demonstrated slightly lower values (0.025 0.120 vs. 0.075 0.124). However, posterior areal strain trended higher in the trivial to mild TR cohort compared to the mild or greater cohort (0.302 0.381 vs. 0.106 0.282). This suggests that patients in the trivial and mild cohort may have more extensible posterior leaflets. Consistent with the areal strain findings, the trivial to mild TR cohort exhibited lower average 1 principal strain in the anterior leaflet compared to the mild or greater cohort (0.303 0.113 vs. 0.410 0.158), but higher average 1 principal strain in the posterior leaflet (0.517 0.221 vs. 0.463 0.203). For the septal leaflet, the average 1 principal strain was 0.273 0.026 in the trivial to mild cohort and 0.361 0.206 in the mild or greater cohort.
@@ 3.4.3 Population study of strain patterns across the patient cohort
Principal component analyses (PCA) of the TVs at mid-diastole and mid-systole were performed independently to determine the mean valve shape and its standard deviations (SD) across the 11 patients. Leaflet strains were computed by tracking point-to-point correspondences between the open and closed valve geometries using our distance-weighted coherent point drift (DW-CPD) feature tracking approach. The resulting spectrum of valve morphologies and associated strain information is shown in Fig. 4E. Within this spectrum, the valve appeared most tethered at –2 SD from the mean shape and most billowing at +2 SD. Both ends of the morphological spectrum demonstrated markedly elevated strains relative to the mean shape. In the most tethered valve, high strain was distributed across all three leaflets, whereas in the most billowing valve, high strain was concentrated on the septal leaflet.
The mean shape exhibited relatively low strains (–0.114 0.259 average areal strain and 0.148 0.175 average 1 principal strain). In contrast, the most tethered valve showed an average areal strain of 0.169 0.433 and a 1 principal strain of 0.425 0.246, while the most billowing valve demonstrated an average areal strain of –0.203 0.392 and a 1 principal strain of 0.221 0.300. For both areal and 1 principal strain, the SD of the strain distribution increased as valve geometries deviated further from the mean shape. These findings suggest that both elevated 1 principal strain and greater strain variability, represented by SD, can be associated with pathological TV morphology.
3.5 Strain assessment of normal and diseased pediatric mitral valves (N=10)
@@ 3.5.1 Patient population
This study cohort included 10 pediatric patients, aged 2 to 16 years at the time of image acquisition. Of these, 5 had normal coaptation, 3 demonstrated billowing of one or more leaflets, and 2 had restricted posterior leaflets. Further, the cohort included 1 patient with an MV arcade, 1 with rheumatic MV, and 1 with Marfan syndrome. Excellent feature tracking accuracy was obtained, as demonstrated in the visualization of the three cut planes shown. Fig. 5A shows the superimposition of ground truth segmentation and the estimated atrial leaflet surface onto 3D TEE images from two representative patients. The ages and clinical diagnoses of all patients are summarized in Fig. 5B.
@@ 3.5.2 Strain characteristics on individual leaflet
We organized the patient cohort into normal and diseased groups and compared the average areal and 1 principal strains (Fig. 5D). Neither group exhibited notable changes in leaflet area between the mid-diastolic and mid-systolic frames. In the normal cohort, the average areal strain was 0.034 0.147 in the anterior leaflet and 0.029 0.171 in the posterior leaflet. In the diseased cohort, the corresponding values were –0.038 0.198 and 0.018 0.354.
The posterior leaflets exhibited higher 1 principal strain than the anterior leaflets in both normal and diseased cohorts. In the normal cohort, the average 1 principal strain was 0.262 0.152 in the anterior leaflet and 0.337 0.117 in the posterior leaflet. In the diseased cohort, the corresponding values were 0.223 0.124 and 0.388 0.356. No qualitative differences in 1 principal strain were observed between the two cohorts, as the inter-cohort differences for each leaflet were less than 0.05.
@@ 3.5.3 Population study of strain patterns across the patient cohort
Although differences in MV strain were modest, we observed a unidirectional increase in mean areal strain from –0.084 to –0.035 as MV morphology shifted from –2 to +2 SD of valve shapes. In contrast, both ends of the morphological spectrum exhibited higher average 1 principal strain than the mean shape: 0.162 at –2 SD and 0.191 at +2 SD, compared to 0.131 in the mean shape. Similar to the TV analysis, the SD of areal and 1 principal strain trended greater as valve geometries deviated further from the mean.
3.6 Strain assessment of adult mitral valves (N=10)
@@ 3.6.1 Patient population
This study cohort comprised 10 adult cardiac surgery patients, aged 49 to 78 years at the time of intra-operative image acquisition. All patients demonstrated normal MV function with no to mild mitral regurgitation (MR), though each had non-mitral cardiac pathology. Excellent feature tracking accuracy was demonstrated in all patients. The ground truth segmentation and the estimated atrial leaflet surface overlaid onto 3D TEE images from two representative patients are presented in Fig.6A. Further, Fig.6B summarizes the ages and clinical diagnoses of the entire cohort. Unlike the previous two cohorts, none of the patients in this group were identified with billowing leaflets.
@@ 4.1 General comments
In this study, we introduce an open-source, feature tracking-based framework for quantifying leaflet strain in atrioventricular valves. Implementation accuracy was verified using a FEA benchmark problem, where agreement was observed in both shape similarity and strain estimation, with Pearson’s () and Spearman’s () in both areal and 1 principal strain. Relative to BCPD [14] and classic CPD [28], our method consistently achieved excellent performance across all evaluated metrics. To assess robustness to morphological variation, we applied the framework to three patient cohorts: (1) TVs in patients with HLHS, (2) pediatric MVs, and (3) adult MVs. The dataset spans normal to regurgitant valves and ages from neonates to the late seventies. This constitutes the first human analysis of tricuspid valve strain quantification comparing healthy and diseased cohorts, and the most heterogeneous patient population studied to date. Despite substantial variation in valve geometry across patients and cohorts, the same feature tracking parameters produced excellent shape agreement in all cases. This highlights the generalizability of the framework. Across cohorts, the median and IQR of the 1 principal strain distributions emerged as a potential marker of altered leaflet biomechanics, with elevated median and IQR values in valves with leaflet billow.
@@
To provide a summary of our strain analysis for MVs, qualitative differences in the characteristics of 1 principal strain distributions between pediatric and adult MVs were observed. In pediatric valves, the distribution was broader with a higher IQR, whereas in adults, it was narrower and more peaked. In cohort-level PCA, both groups showed increasing strain IQR as valve geometries deviated from the mean shape. In adults, distinct patterns emerged in which strain in the posterior leaflet progressively increased and became more diffuse with increasing anterior–posterior diameter. In contrast, this pattern was not observed in the pediatric MV cohort. In our pediatric cohort (mean age 8.9 4.9 years), the average principal strain was 0.262 in the anterior leaflet vs. 0.337 in the posterior leaflet in valves with normal morphology. Conversely, the average 1 principal strain reported in our adult cohort (mean age 64.6 9.4 years) was markedly lower, with strains exhibiting 0.081 and 0.231, respectively. These preliminary findings are consistent with the general intuition that leaflet strain reduces with age, attributed to an increase in tissue stiffness and thickness.